Subject: APO 3.5m user's committee meeting, 09/16/02

From: strauss@astro.Princeton.EDU

Submitted: Fri, 20 Sep 2002 11:19:25 -0400 (EDT)

Message number: 609 (previous: 608, next: 610 up: Index)

  Apache Point Observatory 3.5m User's Committee Meeting
		September 16, 2002

Attending: Bruce Gillespie, Mike Shull, Ed Turner, Michael Strauss,
Alan Uomoto, Rene Walterbos, Bruce Balick, Jon Holtzman, Chris Stubbs

Absent: John Morse, Al Harper

 --New DIS throughput measurements
 --Results from summer shutdown 
 --Advanced coatings for M2 or M3? 
 --Status of telescope collimation
 --Instrument reports and problems 
 --Appendix A: Discussion of results of DIS throughput measurements
 --Appendix B: Summary of summer shutdown results

*******************Measurement of DIS throughput*******************

Jon Holtzman has analyzed and collated measurements of the throughput
of DIS from data taken in March 2001, March 2002, May 2002, and
September 2002; the changes these datasets sample are:

    March 2001:  old chips and old coatings on telescope
    March 2002:  new chips, but bad blue coating
    May 2002  :  even newer blue chip, new coating
    Sept 2002 :  same, but after realuminization of primary and
	      washing of tertiary

However, the throughput data has been obtained with various gratings, making
comparison a bit more complicated, especially on the red channel.

A summary plot can be found at:

Jon's summary of conclusions from this can be found below as Appendix
A of these minutes.  Note in particular that the news is overall very
good, but that we don't quite have all the data necessary to make
final claims on, e.g., the relative throughputs of the low and medium
gratings on the blue side.  This is obviously important for those
deciding on grating settings for the observing runs, and is thus of
high priority for the upcoming engineering run this week. 

***************Shutdown accomplishment list:*********************

The summer shutdown ended a night or two late, right at the beginning
of September.  The mountain staff put a huge amount of work over the
six-week shutdown, working 60 hours or more a week. The User's
Committee expresses its gratitude for the tremendously hard work that
everyone put in during the shutdown.  These shutdown activities have
been tremendously useful in making the telescope mechanically more
robust and reliable.

  The full list of shutdown accomplishments may be found below as
Appendix B.  The most important items were:
  -The realuminization of M1 (as the DIS throughput measurements make
clear, this caused a substantial gain in throughput)
  -Work on the dome shutter drives, which were on the verge of failing
  -Rebuilt the altitude drive assembly

Gillespie notes that the list in Appendix B is divided up into several
   -Tasks accomplished;
   -Tasks remaining to be done, but which can be carried out during
the day;
   -Tasks remaining to be done, which will have to be done at the next
major shutdown
   -New problems discovered 

The bottom line is that we came out of the shutdown with telescope
working fine, but there were a few problems with instruments, which we
detail below. 

**********************Advanced Coatings for the mirrors?**************

  Gillespie: As part of the National Ignition Facility effort, LLNL
has developed new mirror coatings with a reflectivity of >95% from 300
nm to 3 microns, with no dips.  In contrast, our current coatings give
a reflectivity (when newly aluminized) of 90% off each of the three
surfaces (primary, secondary, and tertiary), with dips, e.g., at
8000A, approaching 80%.  These new coatings consist of silver over
aluminum, plus proprietary layers of dialetrics.  They are described
at this web site:

  Moreover, these coatings are easy to clean, and suggest never having
to realuminize again.  The coating is apparently easy to remove if
need be.  The folks at LLNL are eager to get this technology out to
the astronomical community, and will thus do these coatings "at cost",
i.e., $5000/week for lab time, and no further costs.  Their facility
can handle optics as large as 1 meter in diameter, and so would be
appropriate for our secondary or tertiary, but not our primary.  (Of
course, removing the secondary or tertiary for this purpose and
shipping them to Livermore would mean at least a two-week downtime).
This is starting to get the interest in the astronomical community,
and it is not clear how long the research effort will continue, so we
may have a fairly narrow window of opportunity to take advantage of
this.  However, this new technology has been tested in only one
astronomical context; one of the surfaces of one of the Keck
spectrographs has been coated with this, so the robustness of
the coating to outside conditions is not yet known.  We're in the
process of getting a witness sample to mount on the telescope, and see
how it ages.  There is also a possibility of coating the NMSU 1-m
telescope at APO.

		Collimation monitoring project?

  Gillespie: During post-shutdown recommissioning, we found some
short-timescale decollimation; every few nights, the collimation
appears to worsen considerably (this seems not to be a gradual drift,
but rather a jump in the system).  This can cause up to a 1/4 arcsec
image degradation in quadrature.  We should use Shack-Hartman to check
and fix the collimation, as often as every night or perhaps once every
few nights.  Taking the relevant data would take less than 1/2 hour;
let's start a program to do so, look at the results, and decide what
is really needed.  Note that these data would need to be taken with a
fully equilibrated telescope; thus it could not happen in evening
twilight, but perhaps in morning twilight. Collecting these data would
not only be important for keeping the telescope collimated and
pointing well, but will be significant for better establishing the
priority of fixing the "top-end" of the telescope.

		Instrumental reports and problems

  During the summer shutdown, the SPICAM cryotiger failed.  The
company that manufactures the cryotiger is being slow in doing the
repairs, but finally shipped it back to the mountain on Friday, so it
should arrive Monday.  We hope to be back on the air by the end of the
week (getting these things back on line can be difficult).  We've also
ordered a new cryotiger, for backup for both SPICAM and DIS (which
uses the same system).

 The echelle accidentally warmed up prior to the shutdown, and some
kind of condensate appears to have formed on a cold optic (not the
chip); this is now fixed after much pumping.

  The Nasmyth guider thermoelectric cooler failed; this was sent to
the company, and we haven't heard from them yet as to when they will
send it back to us.  (Late update: this is being repaired now, and may
have it back at APO as early as the end of the week).  The telescope
has been tracking very well lately, and Strauss reported that he had
to do only very minor tweaks to the pointing during 1/2 hour DIS
exposures recently.

  GRIM is going into a 'missing quadrant mode' (we've seen this
before); multiple power cycles seem to clear this up. 
The mountain staff is also looking at the "tuning" of the analog power
supplies, which have been related to similar problems in the past. 

  Shull: Procurement of NIC-FPS optics is proceeding.  The order for
the chip is in (see last month's minutes).  A brief technical
description of instrument is available on the APO web page:
and more details are at:
Details can be found in the PDF document at the bottom of this web

  Uomoto: JHU/Chicago IR spectrograph.  We have close to a final
optical design.  The overall throughput is up to 30%, and it maybe
even better if we get a really good detector.  The mechanical design
is coming along fine as well, thanks to Steve Smee of SDSS fame.
We'll be using ceramic alumininum overcoated gratings, which have high

  DIS: Jeff Morgan has been spending quite a bit of time on new
optics.  The original glass slug used was a bit too short, which
caused various troubles.  They are still hoping to deliver by the end
of the year. 

  Stubbs: A multi-band imager: The conceptual design is underway: SDSS
griz + possibility of IR extension, simultaneous.  Excellent for
relative photometry, especially for transients.  Note that because
clouds are grey, relative photometry can be done even in cirrussy
conditions.  There are still uncertainties in the optical design:
should we go with reimaging optics?  The science case for u-band
imaging as well is still up in the air (you need substantially longer
exposure times to get to similar signal-to-noise ratios), although
Strauss points out the tremendous amount of u-band science that SDSS
(with the same exposure time in all bands) has enabled. 


  With the departure of Camron Hastings, Bill Ketzeback has been hired
as a new observing specialist; he's been on the job 1 1/2 months. He
has previously worked at Lowell Observatory, the Wyoming IR facility,
and Zygo Optical in Tucson.

  Don York and collaborators had planned to submit an NSF proposal to
raise funds for echelle upgrades.  However, they have decided not to
go ahead: a new improved guider/acquisition camera is in hand (it may
be installed in October, or later in the Fall), and the sensitivity of
the current system is somewhat higher than previously thought.

  Proposals from JHU and Colorado for the NIR spectrograph, and for
NIC-FPS, respectively, did go in, however. 

  There was some discussion of the possibility of obtaining a wide
slit (5") for spectrophotometry.  Alan Uomoto expresses interest in
building such a thing, if money is found.  He points out that he has
made an aperture mask with round apertures in the past, with
spectrophotometry in mind. 

  A number of people in the ARC consortium will be present at the
Seattle AAS meeting.  As in previous AAS meetings, we should use the
opportunity to get together privately to discuss various APO-related
matters face-to-face. 

 Al Harper has taken over from Lew Hobbs as the U. Chicago
representative on this user's committee.  We welcome Al to our group,
and thank Lew for his valuable input over the years. 

 Last month's meeting minutes are approved. 

 Next meeting:  Monday, October 14, 11:30 AM

			Appendix A:
Notes from Jon Holtzman on his analysis of throughput data from DIS. 

   1. We've made good throughput progress in the past 18 months with DIS!
      The DIS blue channel in particular should be very competitive (but
      maybe not so great for all gratings, see below).

   2. Realuminization probably improved the throughput about 30 percent in
      the blue channel. Can't say for sure about the red, because we don't
      have data with the same grating. I don't see huge evidence for a
      strong wavelength dependence like we did last time we realuminized,
      but I haven't looked super closely.

   3. It's possible/likely there are significant throughput
      differences between the gratings.

      a. In the blue, comparing March 2002 to May 2002, one sees
      significantly better performance in May with the low res as opposed
      to the medium res: HOWEVER, there was also a change of chip and change
      of coating (though at 5000 A, the latter should not have made much
      difference, and this would be a pretty significant QE difference between
      the two chips). We have no data with the high res grating.

      b. In the red, there's no direct comparison, but the much better
      throughput after aluminization with the low res as compared to before
      with the medium res strongly suggests that the low res grating has
      better throughput.  Also the observation that the old red chips
      with low res grating outperform the new chips with medium and high
      res gratings supports this conclusion.  Medium and high res can
      be compared between March 02 and May 02 directly, suggesting that
      high is better at some wavelengths, worse at others.

   Additional things I would note:

   1. As you can see from scatter between curves, I don't measure the
      throughput at extremely high accuracy. There are several
      complications.  Working slitless makes it difficult to get good
      accuracy; proper sky subtraction is difficult. If we want better
      measurements, we should consider getting a wide slit (e.g. 5-10
      arcsec). Note that all measurements are made on raw (i.e. unflattened)
      data, which I think is appropriate (note that if one flattens with flats
      taken with the slit, one introduces features that come from variations
      along the slit that do not exist in the slitless data), but can
      introduce scatter from poor sky subtraction if the chips aren't flat;
      this is a real problem with the old red chip, because it definitely
      wasn't extremely flat.

   2. In the red channel, the stellar spectra are tilted fairly significantly
      with respect to pixel rows.

   3. I checked the gains at the different epochs, and they are not
      completely consistent with one another. For example, I measured gain
      of 1.74 for the new red channel in March, but got 1.9 for the same
      in September. The plots were made assuming the latter for all dates
      with the new chips; if you adopt the former, the total throughput
      in the red channel goes down by 9 percent. Uncertainties in the gain
      might be arising from non-Gaussian noise in the data, I need to look
      further into this.

	Appendix B: Accomplishments of the summer shutdown (from Mark

	1. reusable safety rails assembled and installed for the
Intermediate and Observing levels for when the large hatches are

	2. The following work was accomplished on both the right and
left shutters
			1. Installed low temp grease in pillow block bearings

			2. installed low temp oil in right angle gear drives

			3. Installed new drive motors

			4. Installed new brakes

			5. Installed new speed reducers with low temp oil

	3. re-aluminized M1

	4. washed M3

	5. calibrated M1 PMSS load cells

	6. rebuilt altitude drive assembly

	7. repaired mirror covers

	8. Evacuated and cooled GrimII

	9. warmed Echelle CCD and started evacuation

	10. Cleaned mirror cell

	11. checked PMSS bellowframs

	12. leaked checked PMSS air supply

	13. Installed exhaust fans for Cryotigers in I-Level racks

	14. Cleaned up Spicam wiring, removed power transformer

	15. Had DIS slit viewer evaluated and moved power supply to
ceiling of I-level and doubled  up 12 volt wiring  to reduce voltage
drop for TEC.

	16. relocated tertiary PCB2 for ease of maintenance

	17. Installed brackets for new cat eye mask

The following items were planned but not completed due to time and manpower

	1. tertiary wiring clean up at PCB3

	2. Shim mirror cover switches

	3. replace some guide rollers on shutter  (parts arrived late)

The following items are still in work awaiting time and or parts

	1. Install new cat eye mask

	2. Evacuate and cool Spicam (needs new cyrotiger parts that
were found contaminated)

	3. Complete Echelle and Spicam evacuation and cool down

	4. DIS off-telescope arc lamp cal system

The following problems were discovered and need further work/investigation

	1. Na2 Guider does not cool properly

	2. The shutter speed reducers are (and never were) appropriate
for this application and require factory modification.

	3. Mirror covers can not support loads over 200 lbs. without
popping rivets

	4. bolt on altitude drive shaft not removable.

	5. PMSS load cells have large drift

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 609 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to